# Multilevel neural simulation-based inference

### v1

@@ 5.4 Cosmological Simulations
Finally, we consider a simulator from cosmology. We use the CAMELS simulation suite (Villaescusa-Navarro et al., 2021, 2023)—a benchmark dataset for machine learning in astrophysics—which comprises both low- and high-fidelity cosmological dataset. Our task is to infer a standard cosmological target parameter using a 39-dimensional power spectra of cosmology data, see Section C.5 for an example. The low-fidelity simulations are gravity-only N-body simulations, whose physical behaviour is controlled only by the parameter and some Gaussian fluctuations of the initial conditions. The high-fidelity hydrodynamic simulations have additional physics, controlled by an additional five parameters. We include these additional parameters as part of the $\mathcal{U}$ space, constituting a case of partial common random numbers between the low- and high-fidelity simulators, similar to Section 5.3.
@@
Here, the high-fidelity simulations can be orders of magnitude (>×100$\times$ 100> × 100) slower to generate than the low-fidelity ones, making this a representative problem for our method. Assuming we only have access to $n=n_{1}=20$ high-fidelity simulations, we wish to ascertain the improvement in inference accuracy obtained by including 1000 low-fidelity simulations (i.e. $n_{0}=980$) using our ML-NPE method. To that end, we measure the NLPD and empirical coverage of the estimated posteriors for 980 test data. The result in Figure 5 shows that ML-NPE performs better than standard NPE for both the metrics. Standard NPE tends to produce overconfident posteriors, while ML-NPE yields calibrated or underconfident posteriors for most confidence levels. Thus, including low-fidelity samples using our method leads to better inference.
@@ 6 Conclusion
This paper shows how to reduce the cost of SBI using MLMC: by leveraging cheap low-fidelity simulations, our method performs accurate inference without requiring many samples from costlier simulators, which is the computational bottleneck in many applications. Moreover, our method does not have any hyperparameter to tune and it can handle more than two fidelity levels, unlike the method of Krouglova et al. (2025). More broadly, our method can be seen as a way to perform multilevel training of conditional density estimators in scenarios where data from different sources with different accuracy levels needs to be combined.
Our method is also particularly appealing since it is complementary to other compute-efficient SBI methods such as sequential SBI (Greenberg et al., 2019; Papamakarios et al., 2019), cost-aware SBI (Bharti et al., 2025), and NPE with self-consistency loss (Schmitt et al., 2024a), among others. A recent contemporary work (Tatsuoka et al., 2025) also proposed a two-step multi-fidelity SBI method, wherein an NPE network is first trained on low-fidelity data, and then the resulting posterior estimate is used to guide sampling from the high-fidelity simulator. This approach could easily be combined with our method by using MLMC to estimate the loss, making it complementary to us as well.
As our method involves gradient adjustments during optimisation, we observe a minor increase of roughly 15%-20% in training time compared to that of standard SBI methods, as reported in Section C.2. This however is not an issue as training time is usually negligible compared to running costly simulators, which is the setting we are in. Another limitation of the MLMC approach is that it is not applicable in cases where seed matching of the low- and high-fidelity simulators is not possible.
@@ Acknowledgments
The authors are grateful to Tim Sullivan for helpful discussions, and to the authors of Krouglova et al. (2025) for sharing their source code. YH and AB were supported by the Research Council of Finland grant no. 362534. FXB was supported by the EPSRC grant [EP/Y022300/1]. NJ was supported by the ERC-selected UKRI Frontier Research Grant EP/Y03015X/1 and by the Simons Collaboration on Learning the Universe.
@@
The appendix is arranged as follows: In Appendix A, we discuss the implementation of the optimiser for our MLMC loss function. Appendix B contains the proofs of the theoretical results presented in Section 4. Lastly, Appendix C consists of the experimental details and additional results.
Appendix A Optimising the MLMC objective
We can rewrite the objective in 2 as
$\ell_{\text{MLMC}}(\phi)$ $=\frac{1}{n_{0}}\sum_{i=1}^{n_{0}}f_{\phi}^{0}(u_{1:m,i}^{0},\theta_{i}^{0})+\sum_{l=1}^{L}\frac{1}{n_{l}}\sum_{i=1}^{n_{l}}f_{\phi}^{l}(u_{1:m,i}^{l},\theta_{i}^{l})+\sum_{l=0}^{L-1}\frac{1}{n_{l+1}}\sum_{i=1}^{n_{l}}-f_{\phi}^{l}(u_{1:m,i}^{l},\theta_{i}^{l})$
$:=\sum_{l=0}^{L}\bar{f}_{\phi}^{l,+}+\sum_{l=0}^{L-1}\bar{f}_{\phi}^{l,-}$ (4)
where $\bar{f}_{\phi}^{l,+}:=\frac{1}{n_{l}}\sum_{i=1}^{n_{l}}f_{\phi}^{l}(u_{1:m,i}^{l},\theta_{i}^{l})$ and $\bar{f}_{\phi}^{l,-}:=\frac{1}{n_{l+1}}\sum_{i=1}^{n_{l+1}}-f_{\phi}^{l}(u_{1:m,i}^{l},\theta_{i}^{l})$. Since $\bar{f}_{\phi}^{l,+}$ and $\bar{f}_{\phi}^{l,-}$ are MC estimates of $\mathbb{E}[f_{\phi}^{l}]$ and $-\mathbb{E}[f_{\phi}^{l}]$ respectively with different number of samples, the loss function involves $L-1$ conflicting objectives.
When applying standard gradient-based stochastic optimisation methods, e.g., (Kingma and Ba, 2015, Adam), we observe that the optimisation process is initially dominated by $\bar{f}_{\phi}^{l,+}$, likely because it contains one additional component $\bar{f}_{\phi}^{L,+}$. As training progresses, the gradient $\nabla_{\phi}\bar{f}_{\phi}^{l,+}$ diminishes, and $\nabla_{\phi}\bar{f}_{\phi}^{l,-}$ begins to dominate, causing an increase in $\bar{f}_{\phi}^{l,+}$ and a decrease in $\bar{f}_{\phi}^{l,-}$; see Figure 7. This dynamic results in a “tug-of-war” between the two components, ultimately leading to unstable gradients and the loss to diverge. Broadly speaking, the core idea behind MLMC is to approximate the objective using $h_{0}(\phi)=\bar{f}_{\phi}^{0,+}$, with subsequent terms serving as corrections. This extends to the optimisation as well: the gradient is primarily determined by $\nabla_{\phi}h_{0}(\phi)=\nabla\bar{f}_{\phi}^{0,+}$, while the corrections are captured by $\nabla_{\phi}h_{l}(\phi)=\nabla_{\phi}\bar{f}_{\phi}^{l,+}+\nabla_{\phi}\bar{f}_{\phi}^{l-1,-}$. Under seed matching, we expect $f_{\phi}^{l}\approx f_{\phi}^{l-1}$, which implies $\bar{f}_{\phi}^{l,+}\approx-\bar{f}_{\phi}^{l-1,-}$ and $\nabla_{\phi}\bar{f}_{\phi}^{l,+}\approx-\nabla_{\phi}\bar{f}_{\phi}^{l-1,-}$, allowing $\nabla_{\phi}h_{l}(\phi)$ to remain small and act as a correction to $\nabla_{\phi}h_{0}(\phi)$ throughout the optimisation. However, in practice, this behaviour does not hold when standard gradient descent is applied naïvely due to the unbalance in the evolution of the gradient over the training epochs.
To address this, we apply two modifications. First, we rescale the gradients in the correction terms to ensure $\|\bar{f}_{\phi}^{l,+}\|\approx\|\bar{f}_{\phi}^{l-1,-}\|$ for all $l=1,\dots,L$, so that $\nabla_{\phi}h_{l}(\phi)$ remains small and consistent with its intended role. See Figure 6 for the illustration in the case of two levels. Note that for two levels, we have $\ell_{\text{MLMC}}(\phi)=h_{0}(\phi)+\bar{f}_{\phi}^{1,+}+\bar{f}_{\phi}^{1,-}$. Second, we incorporate gradient surgery (Liu et al., 2020) to further facilitate optimisation with conflicting objectives. Specifically, when the dot product $\nabla_{\phi}h_{0}(\phi)\cdot\nabla_{\phi}h_{c}(\phi)<0$, where $\nabla_{\phi}h_{c}(\phi):=\sum_{l=1}^{L}\nabla_{\phi}h_{l}(\phi)$, we project each gradient onto the normal plane of the other to remove the conflicting components. We empirically observe that combining these two steps improves the stability of gradient and optimisation throughout the training. Full algorithm is provided in Algorithm 1.
((a))
((b))
((c))
Figure 6: Illustration of the gradient scaling with two levels. The coloured arrows indicate gradient of the correction term, total gradient, and scaled gradient respectively. (a) Ideal case: ||$\nabla_{\phi}h_{c}(\phi)||$ remains small and works as a correction to $\nabla_{\phi}h_{0}(\phi)$. (b) In the later stage of training, $\nabla_{\phi}h_{0}(\phi)$ and $\nabla_{\phi}\bar{f}^{1,+}_{\phi}$ diminishes and $\nabla_{\phi}\bar{f}^{0,-}_{\phi}$ starts to dominate the optimisation. (c) Gradient scale adjustment: We scale $\nabla_{\phi}\bar{f}^{0,-}_{\phi}$ such that $||\bar{f}^{1,+}_{\phi}||\approx||\bar{f}^{0,-}_{\phi}||$ and $||\nabla_{\phi}h_{c}(\phi)||$ remains small throughout the training as intended.
((a)) Loss without gradient adjustment
((b)) Loss with gradient adjustment
Figure 7: Comparison of training losses with and without gradient adjustment. (a) Without gradient adjustment, the contribution of $\bar{f}_{\phi}^{1,-}$ begins to dominate the loss around epoch 1000. The resulting conflict between gradient components leads to unstable optimisation, as evidenced by strong fluctuations in the loss and eventual divergence of $\bar{f}_{\phi}^{1,-}$. (b) With gradient adjustment, all components contribute more stably, and the overall loss decreases steadily eventually, indicating convergence. Loss functions are shown for the g-and-k experiment with NLE.
Algorithm 1 MLMC gradient adjustment
1:Input $\nabla h_{0}(\phi),\{\nabla\bar{f}_{\phi}^{l,+},\nabla\bar{f}_{\phi}^{l-1,-}\}_{l=1}^{L}$
2:for $l=1$ to $L$ do
3:     $\nabla\bar{f}_{\phi}^{l-1,-}$ $\leftarrow\frac{||\nabla\bar{f}_{\phi}^{l,+}||_{2}}{||\nabla\bar{f}_{\phi}^{l-1,-}||_{2}+\epsilon}\nabla\bar{f}_{\phi}^{l-1,-}$
4:end for
5:$\nabla h_{c}(\phi)\leftarrow\sum_{l=1}^{L}(\nabla\bar{f}_{\phi}^{l,+}+\nabla\bar{f}_{\phi}^{l-1,-})$
6:if $\nabla h_{0}(\phi)\cdot\nabla h_{c}(\phi)<0$ then
7:     ~$\tilde{\nabla}h_{0}(\phi)\leftarrow\nabla h_{0}(\phi)-\frac{\nabla h_{0}(\phi)\cdot\nabla h_{c}(\phi)}{||\nabla h_{c}(\phi)||_{2}^{2}}\nabla h_{c}(\phi)$
8:     ~$\tilde{\nabla}h_{c}(\phi)\leftarrow\nabla h_{c}(\phi)-\frac{\nabla h_{0}(\phi)\cdot\nabla h_{c}(\phi)}{||\nabla h_{0}(\phi)||_{2}^{2}}\nabla h_{0}(\phi)$
9:else~$\tilde{\nabla}h_{0}(\phi)\leftarrow\nabla h_{0}(\phi),\tilde{\nabla}h_{1}(\phi)\leftarrow\nabla h_{1}(\phi)$
10:end if
11:Output ~$\tilde{\nabla}\ell_{\text{MLMC}}(\phi)\leftarrow\tilde{\nabla}h_{0}(\phi)+\tilde{\nabla}h_{c}(\phi)$
Appendix B Proof of theoretical results
B.1 Proof of Theorem 1
@@
For the first step, we use a version of a Poincaré-type inequality for log-concave measures due to (Bobkov, 1999) (note that a simpler statement and additional discussion is provided in Proposition 10.1 (b) of Saumard and Wellner (2014) for the case of strongly log-concave measures). This result shows that for any log-concave measure $\mu$, there exists $K_{\text{Poin}}>0$ such that for any sufficiently regular integrand $f:\mathcal{Z}\rightarrow\mathbb{R}$ where $\mathcal{Z}\subseteq\mathbb{R}^{d_{\mathcal{Z}}}$, $\text{Var}[f(z)]\leq K_{\text{Poin}}\mathbb{E}_{z\sim\mu}\left[\|\nabla f(z)\|_{2}^{2}\right]$. Applying this Poincaré inequality to each term of the learning objective (i.e. to the terms in Section B.1 and 6), we get
@@
To simplify notation, we now introduce the vector-valued function $g^{l}(z)=g^{l}(\theta,u_{1:m})=(G^{l}_{\theta}(u_{1})^{\top},\ldots,G^{l}_{\theta}(u_{m})^{\top},\theta)^{\top}$ so that $q_{\phi}^{\text{NLE}}(G^{l}_{\theta}(u)\,|\,\theta)=q_{\phi}^{\text{NLE}}(g^{l}(\theta,u))=q_{\phi}^{\text{NLE}}(g^{l}(z))$. We will now derive our first bound, which looks at the first term in the MLMC expansion. To do so, we simplify 7 as follows
@@
Here, we have that 9 follows due to the chain rule, 10 follows due to the Cauchy-Schwarz inequality of the 2-norm, 11 follows due to the Cauchy-Schwarz inequality of expectations.
We now bound each term in 11 separately. For the first term, we get
@@
where 12 follows from the fact that $(a+b)^{2}\leq 2a^{2}+2b^{2}$, 13 follows from the fact that the Frobenius norm $\|\cdot\|_{F}$ is an upper-bound on the matrix 2-norm, 14 follows from the definition of the Frobenius norm and from the fact that $\mathbb{E}[\|\nabla_{z}\theta\|_{F}^{4}]=d_{\Theta}^{2}$, 15 follows from the fact that $(\sum_{i=1}^{n}a_{i})^{2}\leq n\sum_{i=1}^{n}a_{i}^{2}$, and 16, 17 and 19 follow from the definitions of the 2-norm, the $L^{4}$-norm and the $W^{1,4}$-norm respectively.
We now move on to bounding the second term in 11. Since we assumed in Assumption (A3) that $\nabla\log q^{\text{NLE}}_{\phi}$ satisfies a linear growth condition, we must have that
@@
Here, 21 uses the growth condition, 22 holds due since $(a+b+c)^{4}\leq 27(a^{4}+b^{4}+c^{4})$, and 23 uses the definition of Sobolev norm and the fact that $\mathbb{E}\left[\|\theta\|_{2}^{4}\right]$ is upper bounded by a constant since the expectation is against $\pi$, which a log-concave distribution and hence all its moments are finite (see Section 5.1. of Saumard and Wellner (2014)). Combining Section B.1, 7, 11, 19, and 23 therefore gives:
@@
We can now derive a similar bound for the other terms (i.e. to simplify 8). We do this by first only considering the norm inside of the expectation:
@@
Here, 25 follows from the chain rule, 26 follows by adding and subtracting the term $\nabla_{z}g^{l}(z)\nabla\log q^{\text{NLE}}_{\phi}(g^{l-1}(z))$ and using the triangle inequality, and 27 follows from the Cauchy-Schwarz inequality of the 2-norm. Squaring both sides of this inequality and taking expectations, we then obtain:
@@
where 28 follows from that fact that for $a,b\in\mathbb{R}$, we have $(a+b)^{2}\leq 2a^{2}+2b^{2}$, and 29 follows from the Cauchy-Schwarz inequality for expectations. To conclude this proof, we notice that the derivation from 13 to 19 can be modified by replacing $G^{0}$ by $G^{l}$ to get:
@@
Additionally, we could replace $G^{0}$ by $G^{l-1}$ in the the bound from 21 to 23 in order to get:
@@
×$\qquad\quad\times\mathbb{P}\left[\left\|G^{l}(z)-G^{l-1}(z)\right\|_{2}\geq\delta\right]$
@@
×$\qquad\quad\times\mathbb{P}\left[\left\|G^{l}(z)-G^{l-1}(z)\right\|_{2}<\delta\right]$ (35)
@@
where 35 follows due to the law of total expectation. For 36, the bound on the first term follows due to the triangle inequality and Markov’s inequality, and the bound on the second term follows due to the local-Lipschitz condition (i.e. Assumption (A3)) and the fact that a probability is always upper bounded by $1$. Finally, 37 follows using 33 and the definition of Sobolev norm, and by grouping all constants. Combining all of the above (i.e. 6, 8, 27, 30, 31, 33, and 37), we end up with

### v2

@@ 5.4 Cosmological Simulations
We now consider a cosmological simulator using the CAMELS suite (Villaescusa-Navarro et al., 2021, 2023)—one of the most computationally intensive cosmological simulations to date—which comprises both low and high fidelity data. These are real state-of-the-art simulations being used with real-world observational data. Our task is to infer a standard cosmological target parameter using a 39-dimensional power spectra of cosmology data, see Section˜B.5 for an example. The low-fidelity simulations are gravity-only N-body simulations, whose physical behaviour is controlled only by the parameter and some Gaussian fluctuations of the initial conditions. The high-fidelity hydrodynamic simulations have additional physics, controlled by an additional five parameters. We include these additional parameters as part of the $\mathcal{U}$ space, constituting a case of partial common random numbers between the low- and high-fidelity simulators, similar to Section˜5.3.
@@
Here, the high-fidelity simulations can be orders of magnitude ($\times 100$) slower to generate than the low-fidelity ones, making this a representative problem for our method. Assuming we only have access to $n=n_{1}=20$ high-fidelity simulations, we wish to ascertain the improvement in inference accuracy by including 1000 low-fidelity simulations (i.e. $n_{0}=980$) using our ML-NPE method. To that end, we measure the NLPD and empirical coverage of the estimated posteriors for 980 test data. The result in Figure˜5 shows that ML-NPE performs better than standard NPE for both the metrics. Standard NPE tends to produce overconfident posteriors, while ML-NPE yields calibrated or underconfident posteriors for most confidence levels. Thus, including low-fidelity samples using our method leads to better inference outcomes. Before concluding, we note that several papers demonstrating the potential of multi-fidelity SBI methods in cosmology appeared around the same time as our paper (Saoulis et al., 2025; Thiele et al., 2025). These more in-depth studies clearly highlight the potential for impact of advanced multi-fidelity SBI methods.
@@ 6 Conclusion
This paper demonstrated how to reduce the cost of SBI using MLMC, but could more broadly be seen as a way to perform multilevel training of conditional density estimators in scenarios where data from different sources with different accuracy levels needs to be combined. Our method achieves this without any additional fidelity hyperparameter to tune and can handle more than two fidelity levels. It is also particularly appealing since it is complementary to other compute-efficient SBI methods. For example, Tatsuoka et al. (2025) recently proposed to train an NPE network on low-fidelity data, and to then use the resulting posterior approximation to guide sampling from the high-fidelity simulator. This approach could easily be combined with our method.
In terms of limitations, our method involves gradient adjustments during optimisation, and we did observe a minor increase of roughly 15%-20% in training time compared to that of standard SBI methods (see Section˜B.2). However, this is not a significant issue as training time is usually negligible compared to costly high-fidelity simulations. Another limitation of our approach is that it is not applicable in cases where seed-matching of the low- and high-fidelity simulators is not possible. This was not an issue in any of the examples we encountered, but limit its applicability in some cases.
@@ Acknowledgments
The authors are grateful to Tim Sullivan and David Warne for helpful discussions, and to the authors of Krouglova et al. (2025) for sharing their source code. YH and AB were supported by the Research Council of Finland grant no. 362534. FXB was supported by the EPSRC grant [EP/Y022300/1]. NJ was supported by the ERC-selected UKRI Frontier Research Grant EP/Y03015X/1 and by the Simons Collaboration on Learning the Universe.
@@
The appendix is arranged as follows: Appendix˜A contains the proofs of the theoretical results presented in Section˜4. Appendix˜B consists of the experimental details and additional results.
Appendix A Proof of theoretical results
A.1 Proof of Theorem˜1
@@
For the first step, we use a version of a Poincaré-type inequality for log-concave measures due to [Bobkov, 1999] (note that a simpler statement and additional discussion is provided in Proposition 10.1 (b) of Saumard and Wellner [2014] for the case of strongly log-concave measures). This result shows that for any log-concave measure $\mu$, there exists $K_{\text{Poin}}>0$ such that for any sufficiently regular integrand $f:\mathcal{Z}\rightarrow\mathbb{R}$ where $\mathcal{Z}\subseteq\mathbb{R}^{d_{\mathcal{Z}}}$, $\text{Var}[f(z)]\leq K_{\text{Poin}}\mathbb{E}_{z\sim\mu}\left[\|\nabla f(z)\|_{2}^{2}\right]$. Applying this Poincaré inequality to each term of the learning objective (i.e. to the terms in Section˜A.1 and ˜7), we get
@@
To simplify notation, we now introduce the vector-valued function $g^{l}(z)=g^{l}(\theta,u_{1:m})=(G^{l}_{\theta}(u_{1})^{\top},\ldots,G^{l}_{\theta}(u_{m})^{\top},\theta)^{\top}$ so that $q_{\phi}^{\text{NLE}}(G^{l}_{\theta}(u)\,|\,\theta)=q_{\phi}^{\text{NLE}}(g^{l}(\theta,u))=q_{\phi}^{\text{NLE}}(g^{l}(z))$. We will now derive our first bound, which looks at the first term in the MLMC expansion. To do so, we simplify ˜8 as follows
@@
Here, we have that ˜10 follows due to the chain rule, ˜11 follows due to the Cauchy-Schwarz inequality of the 2-norm, ˜12 follows due to the Cauchy-Schwarz inequality of expectations.
We now bound each term in ˜12 separately. For the first term, we get
@@
where ˜13 follows from the fact that $(a+b)^{2}\leq 2a^{2}+2b^{2}$, ˜14 follows from the fact that the Frobenius norm $\|\cdot\|_{F}$ is an upper-bound on the matrix 2-norm, ˜15 follows from the definition of the Frobenius norm and from the fact that $\mathbb{E}[\|\nabla_{z}\theta\|_{F}^{4}]=d_{\Theta}^{2}$, ˜16 follows from the fact that $(\sum_{i=1}^{n}a_{i})^{2}\leq n\sum_{i=1}^{n}a_{i}^{2}$, and ˜17, ˜18 and ˜20 follow from the definitions of the 2-norm, the $L^{4}$-norm and the $W^{1,4}$-norm respectively.
We now move on to bounding the second term in ˜12. Since we assumed in Assumption (A3) that $\nabla\log q^{\text{NLE}}_{\phi}$ satisfies a linear growth condition, we must have that
@@
Here, ˜22 uses the growth condition, ˜23 holds due since $(a+b+c)^{4}\leq 27(a^{4}+b^{4}+c^{4})$, and ˜24 uses the definition of Sobolev norm and the fact that $\mathbb{E}\left[\|\theta\|_{2}^{4}\right]$ is upper bounded by a constant since the expectation is against $\pi$, which a log-concave distribution and hence all its moments are finite (see Section 5.1. of Saumard and Wellner [2014]). Combining Section˜A.1, ˜8, ˜12, ˜20, and ˜24 therefore gives:
@@
We can now derive a similar bound for the other terms (i.e. to simplify ˜9). We do this by first only considering the norm inside of the expectation:
@@
Here, ˜26 follows from the chain rule, ˜27 follows by adding and subtracting the term $\nabla_{z}g^{l}(z)\nabla\log q^{\text{NLE}}_{\phi}(g^{l-1}(z))$ and using the triangle inequality, and ˜28 follows from the Cauchy-Schwarz inequality of the 2-norm. Squaring both sides of this inequality and taking expectations, we then obtain:
@@
where ˜29 follows from that fact that for $a,b\in\mathbb{R}$, we have $(a+b)^{2}\leq 2a^{2}+2b^{2}$, and ˜30 follows from the Cauchy-Schwarz inequality for expectations. To conclude this proof, we notice that the derivation from ˜14 to ˜20 can be modified by replacing $G^{0}$ by $G^{l}$ to get:
@@
Additionally, we could replace $G^{0}$ by $G^{l-1}$ in the the bound from ˜22 to ˜24 in order to get:
@@
$\qquad\quad\times\mathbb{P}\left[\left\|G^{l}(z)-G^{l-1}(z)\right\|_{2}\geq\delta\right]$
@@
$\qquad\quad\times\mathbb{P}\left[\left\|G^{l}(z)-G^{l-1}(z)\right\|_{2}<\delta\right]$ (36)
@@
where ˜36 follows due to the law of total expectation. For ˜37, the bound on the first term follows due to the triangle inequality and Markov’s inequality, and the bound on the second term follows due to the local-Lipschitz condition (i.e. Assumption (A3)) and the fact that a probability is always upper bounded by $1$. Finally, ˜38 follows using ˜34 and the definition of Sobolev norm, and by grouping all constants. Combining all of the above (i.e. ˜7, ˜9, ˜28, ˜31, ˜32, ˜34, and ˜38), we end up with